OpenStudy (anonymous):

Find the values of x and y that make these triangles congruent by the HL Theorem. x=2, x=3 x=3, y=2 x=1, y=2 x=2, y=1

OpenStudy (anonymous):

You should have a system of equations: x + 5 = y - x [eq 1] y + 5 = 3y + x [eq 2] solve eq 1 for y y = 2x + 5 Eq 2: 2y = 5 - x substitute this expression for y in eq 2 2(2x + 5) = 5 - x expand, combine like terms, solve for x x = - 1 y = 3

OpenStudy (anonymous):

but there is no y=3 wen x=1 /: